OpenStudy (anonymous):

A jeweller had 3 diamonds and 3 different pendants. Two diamonds could be set in each pendant. In one setting, the two diamonds would have a mass of 6 carats. In the second setting, the mass of two diamonds would be 10 carats. In the third, the mass would be 12 carats. Find the mass of each diamonds.

OpenStudy (anonymous):

2 carats, 4 carats and 8 carats 2 + 4 = 6 2 + 8 = 10 4 + 8 = 12 I just used trial and error to find the masses. Not sure how you might do it otherwise.